. Which of the following pairs of ideals in Z[x] are comaximal? · (3), (5) • (3+x), (x) • (5 + x - x2), (3, x)
Answer #1

3) (5) 2x3=6663 and -1X55-56 (5). u= 6 (-(3) and. 6-5=1 -> y = -56 (5) with one with nity = 1. The ideals (3) and (5) are ComUtv = 1 -1 o Hence, (tn) and ) ce mala 3+u)pm + nau, at -) 3 Pu) tn plu) th 980 = 1 => 3 puit n (pnt90)) = 1 + -> 3 Q.+ 20,utn (5th-n?) Tina - (5th-22) =-5-uth2.6 (5th-2²) 3 2x3 = 6 -n=-nxn G (3,0) -> 6th-n? E (3,n) 1 (3,2) are comanemal (-5-u+r?) +